# Markov kernels in Mathlib’s probability library

### v1

@@ 3.1 Proof of the disintegration theorem
Our hypothesis on $\mathcal{Y}$ is not that its $\sigma$-algebra is finite, but that it is countably generated. But that assumption will allow us to reduce to the finite case. For a countably generated $\sigma$-algebra, there exists a sequence of finer and finer $\sigma$-algebras $\mathcal{F}_{n}$, each generated by a partition $\{B_{n,1},\ldots,B_{n,k_{n}}\}$, whose union generates the $\sigma$-algebra of $\mathcal{Y}$.
For each $n$, we can define a density $d_{\kappa,\nu}^{n}$ as above for the finite $\sigma$-algebra $\mathcal{F}_{n}$, which is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and satisfies the integral condition for all sets in $\mathcal{F}_{n}$. The key to turn that sequence into a density is to use a martingale convergence theorem, which was formalized and added to Mathlib in [YD23]. $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ is a filtration, and for each $x\in\mathcal{X}$ and each measurable set $C$ of $\mathcal{Z}$, the sequence of functions $(d_{\kappa,\nu}^{n}(x,\cdot)(C))_{n\in\mathbb{N}}:\mathbb{N}\to\mathcal{Y}\to\mathbb{R}$ can be shown to be a martingale which is bounded in $L^{1}$.
Therefore, by the martingale convergence theorem, it converges almost everywhere to a limit. That limit is almost everywhere equal to the limit superior of the sequence, which is a jointly measurable function. We define $d_{\kappa,\nu}(x,y)(C)$ to be that limit superior. It is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and since it is almost everywhere the limit of the density process in $L^{1}$, it satisfies the integral condition for all measurable sets $C$ of $\mathcal{Z}$.
@@ 3.2 Conditional distributions
The application of disintegration to random variables gives the notion of conditional distribution. Let $(\Omega,P)$ be a probability space, and let $X:\Omega\to\mathcal{X}$ and $Y:\Omega\to\mathcal{Y}$ be two random variables. We can consider the joint distribution of $(X,Y)$, which is the probability measure $(X,Y)_{*}P$ on $\mathcal{X}\times\mathcal{Y}$ that we denote by $P_{XY}$. The distribution of $X$ is likewise denoted by $P_{X}$. If $\mathcal{Y}$ is a standard Borel space, then by the disintegration theorem there exists a kernel $P_{Y|X}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that $P_{XY}=P_{X}\otimes_{m}P_{Y|X}$. That kernel is called a (regular) conditional probability distribution of $Y$ given $X$. This disintegration is exactly how we define conditional distributions in Mathlib.
@@
In a standard Borel space, we can obtain the conditional distributions for the functions $Y=\mathrm{id}_{\Omega}$ and $X$ the identity function on $\Omega$ but seen as a function from $(\Omega,\mathcal{G})$ to $(\Omega,\mathcal{F})$. This is a kernel from $(\Omega,\mathcal{G})$ to $(\Omega,\mathcal{F})$, which we denote by $P^{\mathcal{G}}$ and is called condExpKernel in Mathlib. Its main property is that for every integrable function $f:\Omega\to\mathbb{R}$, the conditional expectation $P[f|\mathcal{G}]$ is almost everywhere equal to the function $\omega\mapsto P^{\mathcal{G}}(\omega)[f]$.
theorem condExp_ae_eq_integral_condExpKernel [NormedAddCommGroup E] {f : Ω E}
@@
That code uses the notation f =ᵐ[μ] g, which means that $f$ and $g$ are equal $\mu$-almost everywhere.
In general the conditional expectations $P(A|\mathcal{G})$ and $P(B|\mathcal{G})$ for different sets $A$ and $B$ are two functions defined only almost everywhere, and for example we have only almost everywhere the equality $P(A\cup B|\mathcal{G})=P(A|\mathcal{G})+P(B|\mathcal{G})$ for $A,B$ disjoint. We can go around that issue as long as only countably many sets are involved, but having a unique version of the conditional expectation that consolidates into a kernel (condExpKernel) is much more convenient since it allows the use of true equality.
@@ 3.3 Posterior kernel
That kernel is called the posterior kernel or Bayesian inverse of $\kappa$ with respect to $\mu$. It is a central object in Bayesian statistics, where $\mu$ is interpreted as the prior distribution on a parameter, $\kappa$ gives the distribution of the observed data given the parameter, and $\kappa_{\mu}^{\dagger}$ maps data to a distribution on the parameter called posterior distribution, that is interpreted as the updated belief about the parameter after observing the data.
A remark about notation: the notation $\kappa_{\mu}^{\dagger}$ makes the measure explicit, which departs from the notation $\kappa^{\dagger}$ used in parts of the literature on posterior kernels [Cle+17], in which the measure is understood from the context. In the code, we have to keep the measure explicit because it cannot be inferred otherwise, and we chose to keep the notation of this paper close to the code.
@@
That lemma states that for $\kappa\circ_{m}\mu$-almost all $y\in\mathcal{Y}$ (written ∀ᵐ y ∂(κ ∘ₘ μ) in code), $\kappa_{\mu}^{\dagger}(y)$ is the measure on $\mathcal{X}$ with density $\omega\mapsto\frac{d\kappa}{d(\kappa\circ_{m}\mu)}(x,y)$ with respect to $\mu$, in which $\frac{d\kappa}{d(\kappa\circ_{m}\mu)}$ is the kernel Radon-Nikodym derivative with respect to the constant kernel with value $\kappa\circ_{m}\mu$. It requires the assumption that there exists a dominating measure: for $\mu$-almost all $x\in\mathcal{X}$, $\kappa(x)$ is absolutely continuous with respect to $\kappa\circ_{m}\mu$. This is in particular always true in the case of a countable space $\mathcal{X}$, for which the result simplifies to
@@ 4 Independence
Those two random variables are conditionally independent given a sub-$\sigma$-algebra $\mathcal{G}$ of the $\sigma$-algebra of $\Omega$ if for all measurable sets $A$ of $\mathcal{X}$ and $B$ of $\mathcal{Y}$, for $\mu$-almost every $\omega\in\Omega$,
@@
Nonetheless, conditional independence is overwhelmingly used in cases where the conditional expectation kernel exists, often to obtain properties of conditional distributions: for example if $X$ is independent of $Y$ given $Z$, the conditional distribution of $X$ given $(Y,Z)$ is almost surely equal to the conditional distribution of $X$ given $Z$ (see further down). In [For21], Patrick Forré introduces a generalization of conditional independence. It is defined for any measurable spaces but he makes the choice of requiring the existence of a conditional distribution to declare that two random variables are conditionally independent given another. Otherwise the property is not deemed useful enough. It is also notable that manipulating conditional independence is much easier when conditional kernels exists, as explained above in the discussion about the conditional expectation kernel. We thus made the choice of accepting the loss of generality for convenience: that choice could be revisited if the need for more general spaces arises. Another more minor refactor would be to follow the example of [For21] and (somewhat artificially) extend our definition to any space by requiring that there exists a conditional kernel such that the definition applies.
The proofs of the properties of independence with respect to a kernel and a measure are thus slightly simplified versions of those of conditional independence. The abstraction of the conditional expectation into a kernel is also surprisingly helpful in practice. First, conditional expectations lead to long expressions which can be hard to read in the infoview of Lean, and the kernel expression is often more readable. But besides that minor convenience, the main advantage is that the rigidity of kernels (due to the different types) prevents us from making mistakes when composing them. Proofs involving conditional independence would involve several $\sigma$-algebras on the same space, hence a function on a space could be measurable with respect to one $\sigma$-algebra but not another, and it is easy to inadvertently mistake one for the other when attempting to code a proof. When using kernels, the different types ensure that we cannot compose two kernels unless the types match precisely, and there is only one $\sigma$-algebra involved on each type. It is often the case in a proof that if two expressions involving kernels have the same type, then they are equal, which makes writing the proofs easier. That is of course not a mathematical truth but only a practical observation: only deterministic kernels satisfy $(\kappa\parallel_{k}\kappa)\circ_{k}\Delta_{\mathcal{X}}=\Delta_{\mathcal{Y}}\circ_{k}\kappa$, although they both have type $\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Y}$.
@@
Our definition of independence with respect to a kernel and a measure is not an attempt to write the most general possible notion of independence, but merely a minimal abstraction of independence and conditional independence (in standard Borel spaces) that allows us to prove the properties we want only once, with a convenient framework. Many generalizations exist and we mention only two here, which also focus on kernels. [For21] defines transitional conditional independence, which is independence of two kernels given another kernel, in a transitional probability space (which is defined by a fourth kernel). We remark that our definition takes inspiration from this and features four objects as well: two random variables, a kernel and a measure. Both random variables and measures are special cases of kernels, and Forré extended independence to the general case of four kernels. [CJ19, Fri20] consider categorical definitions of independence of morphisms in Markov categories (and Markov kernels are such a category).
@@
CondIndepFun (mγ.comap k) hk.comap_le Y X μ
@@
The definition CondIndepFun m hm Y X μ means that $Y$ and $X$ are conditionally independent given the sub-$\sigma$-algebra $m$ of the $\sigma$-algebra of $\Omega$, for the measure $\mu$. hm is a proof that $m$ is indeed a sub-$\sigma$-algebra of the $\sigma$-algebra of $\Omega$. Here the $\sigma$-algebra used is $\sigma(Z)$, the $\sigma$-algebra generated by $Z$, which is obtained by MeasurableSpace.comap Z inferInstance. It is a sub-$\sigma$-algebra because $Z$ is measurable, which is proved by hZ.comap_le. Writing about conditional independence given a random variable $Z$ instead of a sub-$\sigma$-algebra $\mathcal{G}$ is currently rather verbose, and we could consider adding a more convenient notation for that common case in the future. The definition Kernel.prodMkRight turns a kernel $\mathcal{Z}\rightsquigarrow\mathcal{X}$ into a kernel $\mathcal{Z}\times\mathcal{Y}\rightsquigarrow\mathcal{X}$ by ignoring the second argument.
@@ 5 Sub-Gaussian random variables
A sub-Gaussian random variable is a real-valued random variable whose tail decreases at least as fast as a Gaussian. We refer the reader to [Ver18] for a thorough introduction to sub-Gaussian random variables and their properties. Formally, $X:\Omega\to\mathbb{R}$ is sub-Gaussian with constant $\sigma^{2}$ on a probability space $(\Omega,P)$ if $e^{uX}$ is integrable for all $u\in\mathbb{R}$ and the function $u\mapsto P\left[e^{uX}\right]$ (its moment generating function, MGF) satisfies the inequality $P\left[e^{uX}\right]\leq e^{u^{2}\sigma^{2}/2}$ for all $u\in\mathbb{R}$. The function $u\mapsto e^{u^{2}\sigma^{2}/2}$ is the MGF of a real Gaussian random variable with mean 0 and variance $\sigma^{2}$, hence the name sub-Gaussian. Sub-Gaussianity is a useful property to prove concentration inequalities such as Hoeffding’s inequality [Hoe63], which states in the independent case that if $X_{1},\ldots,X_{n}$ are independent and sub-Gaussian with constant $\sigma^{2}$, then for all $t>0$,
@@
The := by volume_tac code defines a default value in case no measure is provided. That default value uses a tactic (volume_tac) that infers the measure if $\Omega$ has been endowed with a MeasureSpace instance. For example if $\Omega=\mathbb{R}$, we don’t need to specify that we mean sub-Gaussian with respect to the Lebesgue measure.
The conditional version of the sub-Gaussian property is defined similarly, by replacing the expectation $P[\cdot]$ by the conditional expectation $P[\cdot|\mathcal{G}]$ given a sub-$\sigma$-algebra $\mathcal{G}$ of the $\sigma$-algebra of $\Omega$. A random variable $X:\Omega\to\mathbb{R}$ is conditionally sub-Gaussian with constant $\sigma^{2}$ given $\mathcal{G}$ if $e^{uX}$ is integrable for all $u\in\mathbb{R}$ and $P$-almost surely, $P\left[e^{uX}\mid\mathcal{G}\right]\leq e^{u^{2}\sigma^{2}/2}$ for all $u\in\mathbb{R}$. We note that this is almost never mentioned as a separate definition in textbooks, which implicitly assume that conditional versions of properties are defined by replacing expectations by conditional expectations. That definition is not however how we defined it in Mathlib: we instead introduce a new notion of sub-Gaussian with respect to a kernel and a measure, and define conditionally sub-Gaussian as a particular case of that definition for the conditional expectation kernel.
@@
The integrability condition in Definition 5.1 merits discussion. In the unconditional case, we require that $\exp(uX)$ is integrable with respect to $P$, which could be interpreted as a technical condition to ensure that the MGF that appears in the inequality is well-defined. The analogue of that for the kernel case would be to require that for $\mu$-almost all $t\in\mathcal{T}$, for all $u$, $\exp(uX)$ is integrable with respect to $\kappa(t)$ (and this is how we implemented it originally). This is implied by the integrability with respect to $\kappa\circ_{m}\mu$, but is a weaker condition. Let’s call it weak integrability. Integrability of a measurable function $f$ with respect to $\kappa\circ_{m}\mu$ is equivalent to $\mu$-almost everywhere integrability of $f$ with respect to $\kappa(t)$ together with integrability of $t\mapsto\int_{y}\|f(y)\|d\kappa(t)$. The weak integrability condition that ensures only that the MGF inequality makes sense has a flaw: it does not compose well.
@@
That lemma is the right tool to prove concentration inequalities for sums of sub-Gaussian random variables in the Ionescu-Tulcea setting of a sequence of kernels. In order to prove the MGF inequality in that result, we have to obtain at least that for $\nu$-almost all $t\in\mathcal{T}$, for all $u\in\mathbb{R}$, $\exp(u(X+Y))$ is integrable with respect to $(\kappa\otimes_{k}\eta)(t)=\kappa(t)\otimes_{m}\eta$, in order for the MGF to be defined. With the weak integrability condition, we only know that for $\nu$-almost all $t\in\mathcal{T}$, for all $u$, $\exp(uX)$ is integrable with respect to $\kappa(t)$ and that for $(\nu\otimes_{m}\kappa)$-almost all $(t,x)\in\mathcal{T}\times\mathcal{X}$, for all $u$, $\exp(uY)$ is integrable with respect to $\eta(t,x)$. That is not enough to conclude, mainly because the result of integrability with respect to $\kappa(t)\otimes_{m}\eta$ is stronger than $\kappa(t)$-almost everywhere integrability with respect to $\eta(t,x)$. The stronger integrability condition in Definition 5.1 solves that issue.

### v2

@@ 3.1 Proof of the disintegration theorem
Our hypothesis on $\mathcal{Y}$ is not that its $\sigma$-algebra is finite, but that it is countably generated. But that assumption will allow us to reduce to the finite case. For a countably generated $\sigma$-algebra, there exists a sequence of finer and finer $\sigma$-algebras $\mathcal{F}_{n}$, each generated by a partition $\{B_{n,1},\ldots,B_{n,k_{n}}\}$, whose union generates the $\sigma$-algebra of $\mathcal{Y}$ (see countablePartition for the partition into finitely many sets, and countableFiltration for the filtration built out of them).
For each $n$, we can define a density $d_{\kappa,\nu}^{n}$ as above for the finite $\sigma$-algebra $\mathcal{F}_{n}$, which is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and satisfies the integral condition for all sets in $\mathcal{F}_{n}$.
[…]
x (countablePartitionSet n y ×ˢ C) / ν x (countablePartitionSet n y)).toReal
Here countablePartitionSet n y is the set in the partition $\{B_{n,1},\ldots,B_{n,k_{n}}\}$ containing $y$. The key to turn that sequence into a density is to use a martingale convergence theorem, which was formalized and added to Mathlib by [ying2023formalization]. $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ is a filtration, and for each $x\in\mathcal{X}$ and each measurable set $C$ of $\mathcal{Z}$, the sequence of functions $(d_{\kappa,\nu}^{n}(x,\cdot)(C))_{n\in\mathbb{N}}:\mathbb{N}\to\mathcal{Y}\to\mathbb{R}$ can be shown to be a martingale which is bounded in $L^{1}$.
lemma martingale_densityProcess (hκν : fst κ ν) [IsFiniteKernel ν]
[…]
Therefore, by the martingale convergence theorem, it converges almost everywhere to a limit (Submartingale.tendsto_eLpNorm_one_limitProcess). That limit is almost everywhere equal to the limit superior of the sequence, which is a jointly measurable function. We define $d_{\kappa,\nu}(x,y)(C)$ to be that limit superior.
[…]
limsup (fun n densityProcess κ ν n x y C) atTop
It is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and since it is almost everywhere the limit of the density process in $L^{1}$, it satisfies the integral condition for all measurable sets $C$ of $\mathcal{Z}$.
@@ 3.2 Conditional distributions
The application of disintegration to random variables gives the notion of conditional distribution. Let $(\Omega,P)$ be a probability space, and let $X:\Omega\to\mathcal{X}$ and $Y:\Omega\to\mathcal{Y}$ be two random variables. We can consider the joint distribution of $(X,Y)$, which is the probability measure $(X,Y)_{*}P$ on $\mathcal{X}\times\mathcal{Y}$ (pushforward of $P$ by the map $(X,Y)$) that we denote by $P_{XY}$. The distribution of $X$ is likewise denoted by $P_{X}$. If $\mathcal{Y}$ is a standard Borel space, then by the disintegration theorem there exists a kernel $P_{Y|X}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that $P_{XY}=P_{X}\otimes_{m}P_{Y|X}$. That kernel is called a (regular) conditional probability distribution of $Y$ given $X$. This disintegration is exactly how we define conditional distributions in Mathlib (condDistrib).
@@
In a standard Borel space, we can obtain the conditional distributions for the functions $Y=\mathrm{id}_{\Omega}$ (from $(\Omega,\mathcal{F})$ to itself) and $X$ the identity function on $\Omega$ but seen as a function from $(\Omega,\mathcal{F})$ to $(\Omega,\mathcal{G})$. This is a kernel from $(\Omega,\mathcal{G})$ to $(\Omega,\mathcal{F})$, which we denote by $P^{\mathcal{G}}$ and is called condExpKernel in Mathlib. Its main property is that for every integrable function $f:\Omega\to\mathbb{R}$, the conditional expectation $P[f|\mathcal{G}]$ is almost everywhere equal to the function $\omega\mapsto P^{\mathcal{G}}(\omega)[f]$.
theorem condExp_ae_eq_integral_condExpKernel {m mΩ : MeasurableSpace Ω}
: Measure Ω} [IsFiniteMeasure μ] [NormedAddCommGroup E] {f : Ω E}
@@
That code uses the notation f =ᵐ[μ] g, which means that $f$ and $g$ are equal $\mu$-almost everywhere (the character is part of the notation and is not related to the measurable space m).
In general the conditional expectations $P(A|\mathcal{G})$ and $P(B|\mathcal{G})$ for different sets $A$ and $B$ are two functions defined only almost everywhere, and for example we have only almost everywhere the equality $P(A\cup B|\mathcal{G})=P(A|\mathcal{G})+P(B|\mathcal{G})$ for $A,B$ disjoint. We can go around that issue as long as only countably many sets are involved, but having a unique version of the conditional expectation that consolidates into a kernel (condExpKernel) is much more convenient since it allows the use of equality instead of almost-everywhere equality.
@@ 3.3 Posterior kernel
The kernel $\kappa_{\mu}^{\dagger}$ is called the posterior kernel or Bayesian inverse of $\kappa$ with respect to $\mu$. It is a central object in Bayesian statistics, where $\mu$ is interpreted as the prior distribution on a parameter, $\kappa$ gives the distribution of the observed data given the parameter, and $\kappa_{\mu}^{\dagger}$ maps data to a distribution on the parameter called posterior distribution, that is interpreted as the updated belief about the parameter after observing the data.
A remark about notation: the notation $\kappa_{\mu}^{\dagger}$ makes the measure explicit, which departs from the notation $\kappa^{\dagger}$ used in parts of the literature on posterior kernels [clerc2017pointless], in which the measure is understood from the context. In the code, we have to keep the measure explicit because it cannot be inferred otherwise, and we chose to keep the notation of this paper close to the code.
@@
That lemma states that for $\kappa\circ_{m}\mu$-almost all $y\in\mathcal{Y}$ (written ∀ᵐ y ∂(κ ∘ₘ μ) in code), $\kappa_{\mu}^{\dagger}(y)$ is the measure on $\mathcal{X}$ with density $x\mapsto\frac{d\kappa}{dc_{(\kappa\circ_{m}\mu)}}(x,y)$ with respect to $\mu$, in which $c_{(\kappa\circ_{m}\mu)}$ is the constant kernel with value $\kappa\circ_{m}\mu$. It requires the assumption that there exists a dominating measure: for $\mu$-almost all $x\in\mathcal{X}$, $\kappa(x)$ is absolutely continuous with respect to $\kappa\circ_{m}\mu$. This is in particular always true in the case of a countable space $\mathcal{X}$, for which the result simplifies to
@@ 4 Independence
Those two random variables are conditionally independent given a sub-$\sigma$-algebra $\mathcal{G}$ of the $\sigma$-algebra of $\Omega$ if for all measurable sets $A$ of $\mathcal{X}$ and $B$ of $\mathcal{Y}$, for $P$-almost every $\omega\in\Omega$,
@@
Nonetheless, conditional independence is overwhelmingly used in cases where the conditional expectation kernel exists, often to obtain properties of conditional distributions: for example if $X$ is independent of $Y$ given $Z$, the conditional distribution of $X$ given $(Y,Z)$ is almost surely equal to the conditional distribution of $X$ given $Z$ (see further down). [forre2021transitional] introduces a generalization of conditional independence. It is defined for any measurable spaces but he makes the choice of requiring the existence of a conditional distribution to declare that two random variables are conditionally independent given another. Otherwise the property is not deemed useful enough. It is also notable that manipulating conditional independence is much easier when conditional kernels exists, as explained above at the end of Section 3.2. We thus made the choice of accepting the loss of generality for convenience: that choice could be revisited if the need for more general spaces arises. Another more minor refactor would be to follow the example of [forre2021transitional] and extend our definition to any space by requiring that there exists a conditional kernel such that the definition applies.
The proofs of the properties of independence with respect to a kernel and a measure are thus slightly simplified versions of those of conditional independence. The abstraction of the conditional expectation into a kernel is also surprisingly helpful in practice. First, conditional expectations lead to long expressions which can be hard to read in the view of a Lean editor showing the context and goal, and the kernel expression is often more readable. But besides that minor convenience, the main advantage is that the rigidity of kernels (due to the different types) prevents us from making mistakes when composing them. Proofs involving conditional independence would involve several $\sigma$-algebras on the same space, hence a function on a space could be measurable with respect to one $\sigma$-algebra but not another, and it is easy to inadvertently mistake one for the other when attempting to code a proof. When using kernels, the different types ensure that we cannot compose two kernels unless the types match precisely, and there is only one $\sigma$-algebra involved on each type. It is often the case in a proof that if two expressions involving kernels have the same type, then they are equal, which makes writing the proofs easier. That is of course not a mathematical truth but only a practical observation: only deterministic kernels satisfy $(\kappa\parallel_{k}\kappa)\circ_{k}\Delta_{\mathcal{X}}=\Delta_{\mathcal{Y}}\circ_{k}\kappa$, although they both have type $\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Y}$.
@@
Our definition of independence with respect to a kernel and a measure is not an attempt to write the most general possible notion of independence, but merely a minimal abstraction of independence and conditional independence (in standard Borel spaces) that allows us to prove the properties we want only once, with a convenient framework. Many generalizations exist and we mention only two here, which also focus on kernels. [forre2021transitional] defines transitional conditional independence, which is independence of two kernels given another kernel, in a transitional probability space (which is defined by a fourth kernel). We remark that our definition takes inspiration from this and features four objects as well: two random variables, a kernel and a measure. Both random variables and measures are special cases of kernels, and Forré extended independence to the general case of four kernels. [cho2019disintegration, fritz2020synthetic] consider categorical definitions of independence of morphisms in Markov categories (and Markov kernels are such a category).
@@
CondIndepFun (mγ.comap Z) hZ.comap_le Y X μ
@@
The definition CondIndepFun m hm Y X μ means that $Y$ and $X$ are conditionally independent given the sub-$\sigma$-algebra $m$ of the $\sigma$-algebra of $\Omega$, for the measure $\mu$. The argument hm is a proof that $m$ is indeed a sub-$\sigma$-algebra of the $\sigma$-algebra of $\Omega$. Here the $\sigma$-algebra used is $\sigma(Z)$, the $\sigma$-algebra generated by $Z$, which is obtained by MeasurableSpace.comap Z inferInstance. It is a sub-$\sigma$-algebra because $Z$ is measurable, which is proved by hZ.comap_le. Writing about conditional independence given a random variable $Z$ instead of a sub-$\sigma$-algebra $\mathcal{G}$ is currently rather verbose, and we could consider adding a more convenient notation for that common case in the future. The definition Kernel.prodMkRight turns a kernel $\mathcal{Z}\rightsquigarrow\mathcal{X}$ into a kernel $\mathcal{Z}\times\mathcal{Y}\rightsquigarrow\mathcal{X}$ by ignoring the second argument.
@@ 5 Sub-Gaussian random variables
A sub-Gaussian random variable is a real-valued random variable whose tail decreases at least as fast as a Gaussian. We refer the reader to [vershynin2018high] for a thorough introduction to sub-Gaussian random variables and their properties. For $X:\Omega\to\mathbb{R}$ a random variable, the function $u\mapsto P\left[e^{uX}\right]$ is called the moment generating function (MGF) of $X$ (recall that $P[\cdot]$ denotes the expectation under the probability measure $P$). Formally, $X:\Omega\to\mathbb{R}$ is sub-Gaussian with constant $\sigma^{2}$ on a probability space $(\Omega,P)$ if $e^{uX}$ is integrable for all $u\in\mathbb{R}$ and the MGF $u\mapsto P\left[e^{uX}\right]$ satisfies the inequality $P\left[e^{uX}\right]\leq e^{u^{2}\sigma^{2}/2}$ for all $u\in\mathbb{R}$. The function $u\mapsto e^{u^{2}\sigma^{2}/2}$ is the MGF of a real Gaussian random variable with mean 0 and variance $\sigma^{2}$, hence the name sub-Gaussian. Sub-Gaussianity is a useful property to prove concentration inequalities such as Hoeffding’s inequality [hoeffding1963probability], which states in the independent case that if $X_{1},\ldots,X_{n}$ are independent and sub-Gaussian with constant $\sigma^{2}$, then for all $t>0$,
@@
The := by volume_tac code defines a default value in case no measure is provided. That default value uses a tactic (volume_tac) that infers the measure if $\Omega$ has been endowed with a MeasureSpace instance. For example if $\Omega=\mathbb{R}$, we do not need to specify that we mean sub-Gaussian with respect to the Lebesgue measure.
The conditional version of the sub-Gaussian property is defined similarly, by replacing the expectation $P[\cdot]$ by the conditional expectation $P[\cdot|\mathcal{G}]$ given a sub-$\sigma$-algebra $\mathcal{G}$ of the $\sigma$-algebra of $\Omega$. A random variable $X:\Omega\to\mathbb{R}$ is conditionally sub-Gaussian with constant $\sigma^{2}$ given $\mathcal{G}$ if $e^{uX}$ is integrable for all $u\in\mathbb{R}$ and $P$-almost surely, $P\left[e^{uX}\mid\mathcal{G}\right]\leq e^{u^{2}\sigma^{2}/2}$ for all $u\in\mathbb{R}$. We note that this is almost never mentioned as a separate definition in textbooks [boucheron2013concentration, vershynin2018high, lattimore2020bandit], which implicitly assume that conditional versions of properties are defined by replacing expectations by conditional expectations. That definition is not however how we defined it in Mathlib: we instead introduce a new notion of sub-Gaussian with respect to a kernel and a measure, and define conditionally sub-Gaussian as a particular case of that definition for the conditional expectation kernel.
@@
The integrability condition in Definition 5.1
In the unconditional case, we require that $\exp(uX)$ is integrable with respect to $P$, which could be interpreted as a technical condition to ensure that the MGF that appears in the inequality is well-defined. The analogue of that for the kernel case would be to require that for $\mu$-almost all $t\in\mathcal{T}$, for all $u$, $\exp(uX)$ is integrable with respect to $\kappa(t)$ (and this is how we implemented it originally). This is implied by the integrability with respect to $\kappa\circ_{m}\mu$, but is a weaker condition. Let’s call it weak integrability. Integrability of a measurable function $f$ with respect to $\kappa\circ_{m}\mu$ is equivalent to $\mu$-almost everywhere integrability of $f$ with respect to $\kappa(t)$ together with integrability of $t\mapsto\int_{y}\|f(y)\|d\kappa(t)$. The weak integrability condition that ensures only that the MGF inequality makes sense has a flaw: it does not compose well.
@@
That lemma is the right tool to prove concentration inequalities for sums of sub-Gaussian random variables with respect to a sequence of kernels. In order to prove the MGF inequality in that result, we have to obtain at least that for $\nu$-almost all $t\in\mathcal{T}$, for all $u\in\mathbb{R}$, $\exp(u(X+Y))$ is integrable with respect to $(\kappa\otimes_{k}\eta)(t)=\kappa(t)\otimes_{m}\eta$, in order for the MGF to be defined. With the weak integrability condition, we only know that for $\nu$-almost all $t\in\mathcal{T}$, for all $u$, $\exp(uX)$ is integrable with respect to $\kappa(t)$ and that for $(\nu\otimes_{m}\kappa)$-almost all $(t,x)\in\mathcal{T}\times\mathcal{X}$, for all $u$, $\exp(uY)$ is integrable with respect to $\eta(t,x)$. That is not enough to conclude, mainly because the result of integrability with respect to $\kappa(t)\otimes_{m}\eta$ is stronger than $\kappa(t)$-almost everywhere integrability with respect to $\eta(t,x)$. The stronger integrability condition in Definition 5.1 solves that issue.